# Binding Selectivity Analysis from Alchemical Receptor Hopping and Swapping Free Energy Calculations

### v2

@@ 3.1 System Setup and Simulation Settings
The guests and the ligands were docked into the respective receptor binding sites hosts using Maestro (Schrödinger, Inc.). A flat-bottom harmonic positional restraint with a force constant $k_{c}=25$ kcal/mol/Å2 and tolerance of $4.5$ on the centers of mass of the ligand and the receptor as in reference 30 was applied to define the binding site region.32, 33 In the RBFE calculations, alignment restraints24 were applied to each pair of guests using the reference atoms and forced constants listed in the Supplementary Information. Guest and ligands were assigned GAFF2 parameters, and the protein receptors were parameterized using the Amber ff14SB force field.34 The receptors and ligands were combined on tLeaP. The systems were minimized and thermalized before production calculations.
The host-guest alchemical replica exchange molecular dynamics simulations were conducted with the AToM-OpenMM software version 3.5.035 with OpenMM 7.736 and the ATM Meta Force plugin version 0.3.537. The protein-ligand calculations were performed later during the project using the newer 8.1.1 versions of AToM-OpenMM and OpenMM.38, 39 The alchemical calculations employed 22 alchemical states with the non-linear softplus alchemical potential function40, 41 and the alchemical schedules listed in the Supplementary Information. The soft-core perturbation energy function with parameters $u_{\rm max}=200$ kcal/mol, $u_{c}=100$ kcal/mol, and $a=1/16$ was used.41 Molecular dynamics with an MD time-step of 2 fs was conducted for 40 ns per replica for the host-guest calculations and 20 ns per replica for the protein-ligand calculations. Replicas executions were cycled every 40 ps on GPU devices according to the asynchronous replica exchange algorithm.42, 43, 44 Temperature was maintained at 300 K using the Langevin thermostat with a time constant of $0.5$ ps. Perturbation energy samples were collected every 40 ps, and free energies were estimated using the UWHAM method45, 46 after discarding 1/3 of the earliest portion of the trajectory. See https://github.com/Gallicchio-Lab/receptor-hopping to access detailed simulation settings and free energy analysis scripts. Generalized DiffNet analysis was performed using the diffnet-tf package available at https://github.com/Gallicchio-Lab/diffnet-tf.
@@ 4.2 Protein-Ligand Systems
The measured affinities47 translate into a selectivity coefficient of $9.7$ for benzamidine in favor of binding to trypsin over thrombin. In contrast, 1-amidinopipepridine is slightly more selective (a $2.4$ selectivity coefficient) for thrombin over trypsin. Hence, benzamidine is experimentally determined to be $9.7/(1/2.4)\simeq 24$ times more selective for trypsin over thrombin than 1-amidinopipepridine. The corresponding selectivity coefficient ratio calculated from the average of the calculated RSFEs in Table 7 and Eq. (11) is $39$, which is in good agreement with the experiments.
@@
Swapping trypsin with Bz with thrombin with Am $2.05\pm 0.28$ $2.52\pm 0.45$ $1.87$
Swapping of trypsin with Am with thrombin with Bz $-2.32\pm 0.27$ $-2.52\pm 0.45$ $-1.87$
@@ 6 Conclusions
7 Supplementary information
GitHub repository https://github.com/Gallicchio-Lab/receptor-hopping containing the AToM-OpenMM input files for the ABFE, RBFE, RHFE, and RSFE calculations reported in this work.

### v3

@@ 3.1 System Setup and Simulation Settings
The guests and the ligands were docked into the respective receptor binding sites hosts using Maestro (Schrödinger, Inc.). A flat-bottom harmonic positional restraint with a force constant $k_{c}=25$ kcal/mol/Å2 and tolerance of $4.5$ on the centers of mass of the ligand and the receptor as in reference 30 was applied to define the binding site region.32, 33 In the RBFE calculations, alignment restraints24 were applied to each pair of guests using the reference atoms and forced constants as specified in the input files at https://github.com/Gallicchio-Lab/receptor-hopping. Guest and ligands were assigned GAFF2 parameters, and the protein receptors were parameterized using the Amber ff14SB force field.34 The receptors and ligands were combined on tLeaP. The systems were minimized and thermalized before production calculations.
The host-guest alchemical replica exchange molecular dynamics simulations were conducted with the AToM-OpenMM software version 3.5.035 with OpenMM 7.736 and the ATM Meta Force plugin version 0.3.537. The protein-ligand calculations were performed later during the project using the newer 8.1.1 versions of AToM-OpenMM and OpenMM.38, 39 The alchemical calculations employed 22 alchemical states with the non-linear softplus alchemical potential function40, 41 and the alchemical schedules provided in the input files repository at https://github.com/Gallicchio-Lab/receptor-hopping. The soft-core perturbation energy function with parameters $u_{\rm max}=200$ kcal/mol, $u_{c}=100$ kcal/mol, and $a=1/16$ was used.41 Molecular dynamics with an MD time-step of 2 fs was conducted for 40 ns per replica for the host-guest calculations and 20 ns per replica for the protein-ligand calculations. Replicas executions were cycled every 40 ps on GPU devices according to the asynchronous replica exchange algorithm.42, 43, 44 Temperature was maintained at 300 K using the Langevin thermostat with a time constant of $0.5$ ps. Perturbation energy samples were collected every 40 ps, and free energies were estimated using the UWHAM method45, 46 after discarding 1/3 of the earliest portion of the trajectory. See https://github.com/Gallicchio-Lab/receptor-hopping to access detailed simulation settings and free energy analysis scripts. Generalized DiffNet analysis was performed using the diffnet-tf package available at https://github.com/Gallicchio-Lab/diffnet-tf.
@@ 4.2 Protein-Ligand Systems
The measured affinities47 translate into a selectivity coefficient of $9.7$ for benzamidine in favor of binding to trypsin over thrombin. In contrast, 1-amidinopipepridine is slightly more selective (a $2.4$ selectivity coefficient) for thrombin over trypsin. Hence, benzamidine is experimentally determined to be $9.7/(1/2.4)\simeq 24$ times more selective for trypsin over thrombin than 1-amidinopipepridine. The corresponding selectivity coefficient ratio calculated from the average of the calculated RSFEs in Table 7 and Eq. (11) is $81$, which is in reasonably good agreement with the experiments.
@@
Swapping trypsin with Bz with thrombin with Am $2.37\pm 0.30$ $2.52\pm 0.45$ $1.87$
Swapping of trypsin with Am with thrombin with Bz $-2.88\pm 0.30$ $-2.52\pm 0.45$ $-1.87$
@@ 6 Conclusions
7 Acknowledgements
We acknowledge support from the National Institute of General Medical Sciences (NIH 1R15GM151708).
8 Data and Software Availability
The AToM-OpenMM software used in this study is publicly available at https://github.com/Gallicchio-Lab/AToM-OpenMM. GitHub repository https://github.com/Gallicchio-Lab/receptor-hopping contains the AToM-OpenMM input files for the ABFE, RBFE, RHFE, and RSFE calculations reported in this work.